At the beginning of 2017, your company buys a $30,400 piece of equipment that it expects to use for 4 years. The equipment has an estimated residual value of 4,000. The company expects to produce a total of 200,000 units. Actual production is as follows: 50,000 units in 2017, 54,000 units in 2018, 47,000 units in 2019, and 49,000 units in 2020.


Required:

  1. Determine the depreciable cost.
  2. Calculate the depreciation expense per year under the straight-line method.
  3. Use the straight-line method to prepare a depreciation schedule.
  4. Calculate the depreciation rate per unit under the units-of-production method.
  5. Use the units-of-production method to prepare a depreciation schedule.
Answer #1

Answer

· Working

A

Cost

$            30,400.00

B

Residual Value

$              4,000.00

C=A - B

Depreciable base

$            26,400.00

D

Life [in years]

4

E=C/D

Annual SLM depreciation

$              6,600.00

· [a] Depreciable cost = $ 26400

· [b] Depreciation expense per year under Straight line = $ 6600

· [c]

Year

Book Value

Depreciation expense

Ending Book Value

Accumulated Depreciation

2017

$              30,400.00

$              6,600.00

$         23,800.00

$                6,600.00

2018

$              23,800.00

$              6,600.00

$         17,200.00

$             13,200.00

2019

$              17,200.00

$              6,600.00

$         10,600.00

$             19,800.00

2020

$              10,600.00

$              6,600.00

$            4,000.00

$             26,400.00

· [d] Depreciation rate per unit = $ 26400 / 200000 units = $ 0.132

· [e]

Year

Book Value

Usage

Depreciation expense

Ending Book Value

Accumulated Depreciation

1

$              30,400.00

                     50,000

$            6,600.00

$             23,800.00

$           6,600.00

2

$              23,800.00

                     54,000

$            7,128.00

$             16,672.00

$        13,728.00

3

$              16,672.00

                     47,000

$            6,204.00

$             10,468.00

$        19,932.00

4

$              10,468.00

                     49,000

$            6,468.00

$                4,000.00

$        26,400.00